# Remote Work and Women’s Labor Supply: The New Gender Division at Home

### v1

@@ Abstract
We study how increases in remote work opportunities for men affect their spouses’ labor supply. Exploiting variation in the change in work-from-home (WFH) exposure across occupations before and after the COVID-19 pandemic, we find that women whose husbands experienced larger WFH increases are over 2 percentage points more likely to be employed, equivalent to a 4% rise relative to pre-pandemic levels. Evidence from time-use diaries and childcare questionnaires suggests these effects are driven by intra-household reallocation of child-caring time: women are less likely to engage in primary childcare activities, while men working at home partially compensate by covering more for their spouse. These results highlight the role of intra-household spillovers and bargaining in shaping the labor market consequences of remote work.
@@ 1 Introduction
This paper examines whether the expansion of one aspect of workplace flexibility - namely, remote work opportunities - can shift these intra-household dynamics and encourage greater labor force participation among women with children. We focus on U.S. households with young children and combine occupation-level data on remote work vacancies with rich household-level survey data from the CPS and other national datasets. Leveraging quasi-experimental variation in remote work exposure across occupations before and after the pandemic, we estimate whether large-scale changes in remote work availability altered household labor supply decisions, particularly by enabling mothers to remain in or return to the labor force.
Our motivation builds on the idea, emphasized by Goldin (2014), that the last chapter of gender convergence may depend on restructuring work to allow for greater flexibility. Workplace flexibility has been shown to help parents, especially mothers, accommodate both predictable and unexpected caregiving demands (Angelici and Profeta, 2024). Evidence also shows that women and parents of young children place a high value on flexible work arrangements (Mas and Pallais, 2017; Bloom et al., 2022; Aksoy et al., 2022), and are more averse to commuting (Le Barbanchon et al., 2021). While not synonymous with flexibility, remote work is strongly associated with it, particularly in allowing job tasks to be performed from home, thus potentially facilitating a more equitable division of household production.
From the perspective of a collective household model (Chiappori, 1992), increased parental presence at home may ease caregiving constraints and shift bargaining dynamics within couples. Despite this potential, there is limited causal evidence on how remote work affects women’s labor supply, particularly via intra-household channels. Our paper fills this gap by documenting how the expansion of remote work, especially for men, affects their spouses’ employment and time use.
@@
Our first-stage results confirm that the large, occupation-level shifts in remote work feasibility post-COVID led to substantial increases in men’s actual WFH take-up, which almost tripled relative to pre-pandemic levels, with an average rise of about 20 percentage points. In response, we find that wives’ labor market outcomes improved meaningfully: their probability of employment and labor force participation rose by approximately 2.5 percentage points, and their wages increased by around 8%. We also document a 10 percentage point increase in women’s own WFH rates, a rise of roughly 75% relative to their pre-treatment mean. We find some evidence that these positive employment effects for women are mainly driven by households with toddlers but this heterogeneity analysis is not fully robust.
@@
To validate the credibility of our identification strategy, we conduct a broad set of robustness checks that address potential threats to our findings. We examine the possibility of changes in household composition or selection into treatment following the pandemic, including tests for differential fertility and marriage patterns and reweighting strategies that align pre- and post-COVID covariate distributions. We also assess whether our results are driven by specific occupation groups through leave-one-out analyses, and evaluate the sensitivity of our estimates to alternative fixed effect structures and demographic controls. Additionally, we explore the robustness of our findings to varying definitions of the post-treatment period and sample restrictions. Across all these exercises, our main results remain consistent, reinforcing the interpretation that increased WFH options for men played a causal role in raising women’s labor market engagement.
We then proceed to study how shifts in the household’s allocation of labor supply, childcare responsibilities, and household production activities could be driving our employment effects. To do this, we rely on two data sources: the time-use diary data from the American Time Use Survey (ATUS) and the topical modules on childcare from the Survey of Income and Program Participation (SIPP).
@@
Our findings reveal that treated women increase their time spent working, with no significant change in overall household production. However, we observe that women reduce time spent on childcare as a primary activity and husbands timidly seem to compensate for that as both ATUS-based evidence and SIPP topical modules highlight. Households are also less likely to pay for daycare services. We also find evidence that these adjustments in women’s employment and the allocation of childcare are not coming at the expense of the children’s well-being.
@@
Overall, our findings suggest that increases in men’s access to remote work can influence their spouses’ labor supply decisions, particularly in households with young children. Our analysis of time-use and childcare patterns offers tentative evidence that the effects may stem from shifts in household responsibilities: women are reducing their primary childcare time and households spend less in daycare services, while husbands modestly seem to provide more primary childcaring and take care of children while women work. Taken together, the results point to the value of analyzing remote work not only as an individual labor market arrangement, but also as a household-level shock, offering new perspective on how WFH access may affect gender gaps in employment.
@@
Second, we contribute to the growing literature on how working arrangements affect labor market dynamics. While existing research has primarily focused on the direct effects of flexible work arrangements on individual outcomes—highlighting their growing value for both workers and firms (Bloom et al., 2015, 2022; Angelici and Profeta, 2024; Choudhury et al., 2024; Mas and Pallais, 2017)—relatively few studies have examined how changes in work arrangements reshape decision-making at the household level. Prior work shows that women and parents place especially high value on flexibility (Mas and Pallais, 2020; Aksoy et al., 2022), yet evidence on intra-household responses remains limited. We contribute to this emerging strand by leveraging a large-scale, quasi-experimental shift in remote work feasibility to study how increased WFH access for men affects household labor supply and home production. [^1] Closest to our work is Harrington and Kahn (2023), who examine how increases in remote work opportunities, as proxied by women’s college attainment, affect mothers’ employment and earnings trajectories In contrast, we estimate the spillover effects of men’s remote work access on their spouses’ labor market behavior and on the intra-household time allocation.
@@
The remainder of the paper is structured as follows. Section 2 describes our main data sources and outlines the sample restrictions used to construct comparable treatment and control groups. Section 3 presents our empirical strategy and discusses the identification assumptions underlying our analysis. In Section 4, we report the main results on women’s labor market outcomes. Section 5 provides a series of robustness checks and placebo tests to validate our identification strategy. In Section 6, we explore time-use and childcare data to shed light on the mechanisms behind the observed effects. Section 7 concludes.
@@ 2.1 Data Sources
We draw on several public-use U.S. household surveys for our analysis. Our primary results are based on the Current Population Survey (CPS), with complementary evidence from the American Community Survey (ACS) and the Survey of Income and Program Participation (SIPP). To investigate underlying mechanisms, we use detailed time-use information from the American Time Use Survey (ATUS), with additional evidence from SIPP. To measure changes in WFH exposure at the occupation level, we rely on job postings data on remote and hybrid work arrangements (Hansen et al., 2023), as well as established measures of WFH feasibility by occupation from Dingel and Neiman (2020). We describe each data source in greater detail below.
Current Population Survey (CPS)
We use public-use microdata from the CPS (Flood et al., 2024) spanning the years 2016–2024. Our analysis integrates data from both the basic monthly files and the Annual Social and Economic Supplement (ASEC), also known as the March files. The CPS provides detailed information on all household members, including the presence and age of children, as well as occupational codes for both husbands and wives. Critically, it includes several measures of labor supply at the extensive margin, such as employment status and labor force participation, available at both weekly and annual frequencies [^2] See Section 3.3 and Appendix B of Kleven (2024) for a detailed discussion of differences between the CPS basic monthly and March CPS files, and between weekly and annual variables. Weekly variables refer to outcomes in the previous week, while annual variables refer to the previous calendar year., as well as intensive margin measures such part-time status, usual weekly hours worked, and wages [^3] Throughout the paper, we define intensive margin outcomes conditional on employment.. While weekly measures are available in both the basic monthly and March files, annual employment/labor force participation and income variables are available exclusively in the March files. Occupation codes at the 4-digit level are based on the Census Occupation Classification and are already harmonized to the 2010 coding scheme.
WFH Job Postings Data (Hansen et al., 2023).
Our preferred measure of WFH exposure is derived from job postings data on remote and hybrid work arrangements from 2019 to 2024. These data were constructed by a team of researchers in collaboration with Lightcast, using a large language model (LLM) trained to classify job advertisements based on whether the position allows remote work at least one day per week. The resulting measures are available at the month-year–occupation level [^4] Occupations are classified using the Standard Occupational Classification 2018 (SOC) at 3 digits. and at the month-year- industry level. [^5] Industry codes are classified using North American Industry Classification System (NAICS) 2022 at 3 digits. We use this data to build WFH probability changes by occupation, which will be key for our definition of treatment, as we discuss in Subsection 2.3.
@@
American Community Survey (ACS)
We replicate our main findings using ACS (Ruggles et al., 2025) samples from 2017 to 2023. The ACS provides information on current weekly employment status and labor force participation, as well as retrospective measures covering the previous 12 months. [^6] This contrasts with the CPS, where annual outcomes refer to the previous calendar year. These include indicators for employment, months worked, usual hours worked per week, and annual wage and salary income.
Survey of Income and Program Participation (SIPP).
We also replicate our main findings using the 2018-2024 SIPP panel datasets. The SIPP is a nationally representative longitudinal survey that provides comprehensive information for dynamics on household’s employment status and composition, income and government program participation. Starting on 2018, the SIPP has started implementing a 4-year rotational panel survey. [^7] In 2018 the SIPP redesigned the survey that make it difficult to compare the 2014-2017 SIPP panels and post 2018 panels. The 2018-2024 SIPP panels cover information for reference years 2017-2023. Households are interviewed once a year and asked to recall their detailed monthly employment status and income during the previous year. After the year they join the SIPP sample, households are followed and interviewed for the next three years, after which they are dropped from the sample. Each year new households join the sample, so up to 4 different household cohorts populate each year’s SIPP sample. [^8] The 2019 SIPP sample experienced significant logistical difficulties, which led the US Census to prioritize resources on following through with the 2018 sample. As a result, the sample of new households that joined in 2019 was not adequately surveyed and the decision was made to not follow up on them in subsequent surveys. In addition to using the SIPP to validate our cross-sectional findings from the CPS and ACS, we leverage its panel structure to estimate alternative specifications of our main empirical strategy that are only feasible with panel data. [^9] Specifically, we incorporate household fixed effects into our baseline specification to account for time-invariant unobserved heterogeneity. Also, we use SIPP child care topical modules, which are available for all years in our 2018-2024 SIPP panels, to complement mechanism analysis.
Teleworkability Index (Dingel and Neiman, 2020)
As an alternative measure of remote work potential, we incorporate the teleworkability index developed by Dingel and Neiman (2020), which classifies occupations based on their suitability for being performed from home. [^10] The index is constructed using responses from two ONET surveys: the Work Context Questionnaire and the Generalized Work Activities Questionnaire. These surveys assess the importance of physical presence in task performance, allowing for an evaluation of whether specific tasks within each occupation could feasibly be performed remotely. The final occupation-level indicator is calculated as the average teleworkability score across tasks within each occupation. In contrast to our primary measure based on changes in job postings, this index captures the inherent feasibility of telework by occupation, rather than shifts in employer demand for remote work arrangements.

### v2

@@ Abstract
We study how increases in remote work opportunities for men affect their spouses’ labor supply. Exploiting variation in the change in work-from-home (WFH) exposure across occupations before and after the COVID-19 pandemic, we find that increases in husbands’ WFH exposure lead to sizable improvements in wives’ labor-market outcomes: annual employment rises by roughly 2.5 percentage points (from a 69% pre-treatment mean), earnings increase by about 5%, weekly hours worked rise by roughly half an hour, weeks worked increase by about 1.3%, and the likelihood of part-time work declines by approximately 9%. Evidence from time-use diaries and childcare questionnaires suggests these effects are driven by intra-household reallocation of child-caring time: women are less likely to engage in primary childcare activities, while men working at home partially compensate by covering more for their spouse. These results highlight the role of collective household decision-making in shaping the labor market consequences of remote work.
@@ 1 Introduction
This paper examines whether the expansion of one aspect of workplace flexibility - namely, remote work opportunities - can shift these intra-household dynamics and encourage greater labor force participation among women with children. We focus on U.S. households with young children and combine occupation-level data on remote work vacancies with rich household-level survey data from the ACS and other national datasets. Leveraging quasi-experimental variation in remote work exposure across occupations before and after the pandemic, we estimate whether large-scale changes in remote work availability altered household labor supply decisions, particularly by enabling mothers to remain in or return to the labor force.
Our motivation builds on the idea, emphasized by Goldin (2014), that the last chapter of gender convergence may depend on restructuring work to allow for greater flexibility. Workplace flexibility has been shown to help parents, especially mothers, accommodate both predictable and unexpected caregiving demands (Angelici and Profeta, 2024). Evidence also shows that women and parents of young children place a high value on flexible work arrangements (Mas and Pallais, 2017; Bloom et al., 2022; Aksoy et al., 2022), and are more averse to commuting (Le Barbanchon et al., 2021). While not synonymous with flexibility, remote work is strongly associated with it, particularly in allowing job tasks to be performed from home, which could enable a more equitable division of household production between spouses.
We propose a household model of labor supply and childcare with remote work options to guide our empirical analysis. Crucially, the opportunity to look after the child while working poses a technological change in household production function which expands the spouses’ choice set and favors increasing labor supply while making childcare division more equitable. Despite this potential, there is limited causal evidence on how remote work affects women’s labor supply, particularly via intra-household channels. Our paper fills this gap by documenting how the expansion of remote work affects their spousal employment and time use.
@@
Our first-stage estimates indicate that occupation-level increases in WFH probability after COVID translated into substantial increases in men’s realized WFH: take-up rose by about 20 percentage points on average, nearly tripling relative to pre-pandemic levels. In turn, wives’ labor-market outcomes improve along several margins: annual employment rises by roughly 2.5 percentage points (from a 69% pre-treatment mean) or 4%, weekly hours worked rise by roughly half an hour, weeks worked increase by about 1.3%, and the likelihood of part-time work declines by approximately 9%. As a result of this changes in labor supply, overall labor earnings for these women increased by about 5% - with no significant changes in their hourly wage rates. Using a simple decomposition, we show that half of this effect stems from the extensive margin increase and the remainder from the intensive margin.
We then move on to conduct heterogeneity analysis to understand which households have benefited the most. The increases in women’s employment and hours are sharply concentrated among couples with stronger ex ante attachment to remote-capable work: effects are largest for college-educated women (and for women married to college-educated men) and are most pronounced for women ages 25–40, consistent with heightened time constraints during prime childrearing years. Heterogeneity by household composition reinforces a childcare-based channel: responses peak in households with young children—especially when the youngest child is preschool-aged—and attenuate substantially in childless households or those with older children. Finally, we parse out the effect of men’s remote work onto their wives’ labor supply while holding wives’ exposure to remote work fixed through several tests. These show that spousal dynamics and remote work spillovers from one spouse onto the other are economically meaningful, as even women that are unlikely to work from home given their occupation or education degree experienced significant increase in overall labor supply and earnings.
@@
To validate the credibility of our identification strategy, we conduct a broad set of robustness checks that address potential threats to our findings. We examine the possibility of changes in household composition or selection into treatment following the pandemic, including tests for differential fertility and marriage patterns and reweighting strategies that align pre- and post-COVID covariate distributions. We evaluate the sensitivity of our estimates to alternative fixed effect structures, and demographic controls with particular attention to differential trends across major occupational groups, including differences in educational trajectories. Additionally, we explore the robustness of our findings to varying definitions of the post-treatment period and sample restrictions. Across all these exercises, our main results remain consistent, reinforcing the interpretation that increased WFH options for men played a causal role in raising women’s labor market engagement.
@@
Finally, we study how shifts in the household’s allocation of labor supply, childcare responsibilities, and household production activities could be explaining our employment effects. To do this, we rely on mainly on two data sources: the time-use diary data from the American Time Use Survey (ATUS) and the topical modules on childcare from the Survey of Income and Program Participation (SIPP). We complement this data on commuting and remote work from the ACS. Our findings reveal that wives of men that saw large increases in work from home saw a large increase increase their share of day spent working, with no significant change in overall household production. Women themselves are also much more likely to work remote, with approximately a 75% rise relative to the pre -treatment mean, suggesting that remote work enables labor supply for both spouses. Interestingly, we observe that women reduce time spent on childcare as a primary activity and husbands partially with increased childcare while working. Households are also less likely to pay for daycare services. We also find evidence that these adjustments in women’s employment and the allocation of childcare are not coming at the expense of the children’s or women’s well-being.
@@
Overall, our findings suggest that increases in men’s access to remote work can influence their spouses’ labor supply decisions, particularly in households with young children. Our analysis of time-use and childcare patterns offers tentative evidence that the effects may stem from shifts in household responsibilities: women are reducing their primary childcare time and households spend less in daycare services, while husbands modestly seem to provide more primary childcaring and take care of children while women work. Our findings can be rationalized through the lens of a household model where remote work options change the childcare technology and expands households’ feasibility set. Taken together, the results point to the value of analyzing remote work not only as an individual labor market arrangement, but also as a household-level shock, offering new perspective on how WFH access may affect gender gaps in employment.
@@
Second, we contribute to the growing literature on how working arrangements affect labor market dynamics. While existing research has primarily focused on the direct effects of flexible work arrangements on individual outcomes – highlighting their growing value for both workers and firms (Bloom et al., 2015, 2022; Angelici and Profeta, 2024; Choudhury et al., 2024; Mas and Pallais, 2017) – relatively few studies have examined how changes in work arrangements reshape decision-making at the household level. Prior work shows that women and parents place especially high value on flexibility (Mas and Pallais, 2020; Aksoy et al., 2022), yet evidence on intra-household responses remains limited. We contribute to this emerging strand by leveraging a large-scale, quasi-experimental shift in remote work feasibility to study how increased WFH access for men affects household labor supply and home production. [^1] Closest to our work is Harrington and Kahn (2023), who examine how increases in remote work opportunities, as proxied by women’s college attainment, affect mothers’ employment and earnings trajectories In contrast, we estimate the spillover effects of men’s remote work access on their spouses’ labor market behavior and on the intra-household time allocation.
@@
The remainder of the paper is structured as follows. Section 2 presents a simple household model that rationalizes the effects of increasing remote work options on household labor supply and the division of childcare. Section 3 describes our main data sources and outlines the sample restrictions used to construct comparable treatment and control groups. Section 4 presents our empirical strategy and discusses the identification assumptions underlying our analysis. In Section 5, we report the main results on women’s labor market outcomes. Section 6 provides a series of robustness checks and placebo tests to validate our identification strategy. In Section 7, we explore time-use and childcare data to shed light on the intra-household market and non-market supply changes that can help to explain the observed effects on the lens of a collective problem. Section 8 concludes.
2 A Household Model of Remote Work and Labor Supply
To guide the empirical analysis, we present a simple collective household model that rationalizes household decision-making when the option to work remotely becomes available for households with binding childcare constraints. The model formalizes the idea that work from home expands the household’s feasible set of work-care arrangements by allowing some childcare to be produced jointly with market work, thereby relaxing childcare constraints in a collective decision problem.
2.1 Set-up
Consider a dual-parent household composed of a female ($f$), a male ($m$) and a child. Each spouse allocates one unit of time across market work and primary childcare. For the wife, market work is denoted by $h_{f}$ and primary childcare by $n_{f}$. For the husband, total market work is $h_{m}$, which can be performed either on-site ($h_{m}^{O}$) or remotely ($h_{m}^{R}$), and primary childcare is $n_{m}$. Remote work is subject to a technological/feasibility constraint. Specifically, the husband can perform remotely at most a fraction $\bar{r}\in[0,1]$ of his total market work. [^2] We interpret an increase in $\bar{r}$ as an expansion in remote-work opportunities. This is the main source of variation in our empirical analysis. We allow work hours to be bounded above by some idiosyncratic factor $\bar{H}$.
The child in the household is young and their influence in the household comes solely through the demand for childcare. Specifically, the household must meet a minimum childcare time requirement of $\bar{N}$. Primary childcare produced by either spouse contributes one-for-one. In addition, when the husband works remotely, he can provide secondary childcare concurrently with market work at a rate $\bar{\theta}\in[0,1]$. The childcare requirement is thus
$n_{f}+n_{m}+\bar{\theta}h_{m}^{R}\geq\bar{N}.$ (1)
Consumption $c$ is financed by labor earnings from both spouses. Men and women earn wages $w_{m}$ and $w_{f}$, respectively, whether they work from home or not. The consumption good is a numeraire public good. Both spouses have symmetric separable preferences over consumption, market work, and primary childcare:
$U_{i}=u(c)-\varphi(h_{i})-\chi(n_{i}),\qquad i\in\{f,m\},$ (2)
where $u(\cdot)$ is increasing and concave, and $\varphi(\cdot)$ and $\chi(\cdot)$ are increasing and convex. Importantly, disutility from childcare is tied to primary childcare $n_{i}$; secondary childcare generated while working remotely does not involve additional disutility. [^3] This assumption simplifies the analysis and will lead to a 1-to-1 uptake of remote work option. However, as we show in an extension it is not essential to produce our main qualitative predictions. The household solve a standard collective problem, wherein they maximize the weighted sum $\lambda_{f}U_{f}+\lambda_{m}U_{m}$ subject to the aforementioned budget, remote work, childcare and work hours constraints. In what follows, the Lagrange multipliers for these restrictions are $\beta,\mu,\gamma$ and $\alpha_{i}$, respectively.
2.2 Labor and childcare supply
Because primary childcare enters utility with positive marginal disutility, the household has no incentive to supply childcare by any amount higher than $\bar{N}$. Moreover, given our assumptions, remote work weakly dominates on-site work as it generates childcare without incurring in any disutility. The household exhausts available remote work, so $h_{m}^{R}=\bar{r}h_{m}$. An increase in $\bar{r}$ therefore maps one-for-one into higher remote-work hours for the husband, holding total work $h_{m}$ fixed, and mechanically increases secondary childcare $\bar{\theta}h_{m}^{R}$.
The first-order condition with respect to total male labor $h_{m}:=h_{m}^{R}+h_{m}^{O}$ can be written as
$\lambda_{m}\Big[\varphi^{\prime}(h_{m})-\chi^{\prime}(1-h_{m})\Big]-\gamma+\mu\bar{r}+\alpha_{m}+\beta w_{m}=0.$ (3)
An increase in $\bar{r}$ raises the marginal value of male labor through the term $\mu\bar{r}$: additional market work can now be performed remotely which further relaxes childcare constraint. With additional free time, and absent any cap in work hours, the model predicts an increase in total male labor supply. If the male spouse is constrained in work hours, then changes in $\bar{r}$ do not translate into changes in their labor supply, and the primary margin of adjustment is within male labor: the household shifts work towards $h_{m}^{R}$ but keeping $h_{m}$ and $n_{m}$ constant.
From the binding childcare constraint in (1), total primary childcare $n_{f}+n_{m}$ decreases in WFH capacity $\bar{r}$. This time must be reallocated to market work across spouses and the intra-household allocation between spouses is governed by the ratio of marginal utilities, the Pareto weights and the relative wages. However, when the husband is hours-constrained, the adjustment occurs fully through the wife’s primary childcare, substitute $\bar{\theta}\bar{r}\bar{H}$ away from childcare and into their labor supply.
2.3 Extensions
Labor force participation
To capture potential extensive-margin responses, suppose the wife faces a fixed utility cost $F>0$ from entering in the labor market due to, say, commuting/setup costs, job search frictions, or coordination costs of arranging care. Let $V^{work}(\bar{r})$ be the value of the collective problem when the wife works ($h_{f}>0$), and $V^{out}(\bar{r})$ the value when she does not ($h_{f}=0$). When the husband is hours-constrained, increases in $\bar{r}$ raise the marginal benefit of wives’ participation because the household can convert the additional secondary childcare into reduced primary childcare burden only if the wife reallocates time toward work while the husband maintains fixed hours. Moreover, convexities in $\varphi(\cdot)$ and $\chi(\cdot)$ imply gains from diversification across spouses in market and home production. Consequently, the gap $V^{work}(\bar{r})-V^{out}(\bar{r})$ is increasing in $\bar{r}$, making it more likely that the wife enters employment as remote-work opportunities expand.
Remote work for women
Suppose both male and female spouses can work remote a fraction $\bar{r}_{m}$ and $\bar{r}_{f}$ of their overall labor supply, and that their shocks are correlated so that $r_{i}=s_{i}+\rho$. In the absence of any differences in wages or Pareto weights, an idiosyncratic shock on the female’s remote work options, $s_{i}$, raises their labor supply proportionally more than the male. If the shock affects both spouses homogeneously through $\rho$ - say, if both spouses work in the same occupation - then the time gains from remote work that are shifted into labor are split evenly between both spouses. Overall, the effect of remote work shocks on female labor supply increases when both spouses have the option to work remotely.
Remote work for women can also directly change extensive margin of female labor supply if it reduces the fixed cost of women going into the labor market for the first time or after a long unemployment spell. Households will find it less costly to send the female spouse on the labor market if transportation frictions and costs associated with finding and adapting to a new job are significantly reduced. This can be achieved if fixed costs introduced in the previous extension are some decreasing function of the spouse’s common remote work option, $F=F(\rho)$, or are negatively correlated with it.
Demand for daycare
Suppose spouses can hire daycare services $d$ which contributes 1-for-1 towards the childcare restriction at a price of $p$. For rich households that can afford to hire daycare to fully offest primary childcare - i.e. hire $d=\bar{N}$ -, remote work directly reduces their demand for daycare since, if anything, childcare while working substitutes daycare and frees up resources foregone on consumption. [^4] If increasing hours worked is not optimal for the household, then they will reduce demand for daycare by at least the amount that of work pre-shock can be used for childcare, $\bar{\theta}\bar{r}h^{R}_{m}$, and spend it on consumption. For households that cannot afford enough daycare to fully offset daycare - i.e. $d<\bar{N}$ - the overall effect of remote work on daycare demand is ambiguous. Increasing daycare when the technology to do childcare while working is available is inefficient on its own as the household may be providing more childcare than the requirement $\bar{N}$. However, the household may decide to substitute primary childcare for daycare since labor supply (and consequently, earnings) increase following remote work shocks and childcare generates disutility.
Disutility from secondary childcare
The assumption that secondary childcare generates no additional disutility simplifies the main analysis by ensuring full remote work uptake, i.e. $h^{R}_{m}=\bar{r}h_{m}$. We could relax this assumption, for example, by allowing additional secondary childcare to enter the childcare disutility function as $\chi(n_{m}+\bar{\theta}\bar{r}h^{R}_{m})$. Unlike primary childcare, secondary childcare generates labor earnings. Then the overall choice of remote work depends on whether the utility value of the increase in earnings from switching primary childcare into secondary childcare offsets the marginal disutility of secondary childcare. However, the continuity and convexity in the utility functions ensures that households find it optimal to do some positive amount of remote work. [^5] In other words, it cannot be optimal for the household to supply the same amount of labor and childcare without engaging in any remote work as the potential gains in earnings from working a marginal unit of time remotely have not been redistributed across the household. This would only be an option if there was a discontinuous jump in disutility from the first time unit of secondary childcare. Hence the predictions from the baseline model hold qualitatively.
@@ 3.1 Data Sources
We draw on several public-use U.S. household surveys for our analysis. Our primary results are based on the American Community Survey (ACS), with complementary evidence from the Current Population Survey (CPS) and the Survey of Income and Program Participation (SIPP). To investigate underlying mechanisms, we use detailed time-use information from the American Time Use Survey (ATUS), with additional evidence from SIPP. To measure changes in WFH exposure at the occupation level, we rely on job postings data on remote and hybrid work arrangements (Hansen et al., 2023), as well as established measures of WFH feasibility by occupation from Dingel and Neiman (2020). We describe each data source in greater detail below.
American Community Survey (ACS)
We use public-use microdata from ACS (Ruggles et al., 2025) samples from 2017 to 2023. The ACS provides detailed information on all household members, including the presence and age of children, as well as occupational codes for both husbands and wives. [^6] Occupation codes at the 4-digit level are based on the Census Occupation Classification and are already harmonized to the 2010 coding scheme. Critically, the ACS provides information on current weekly employment status and labor force participation, as well as retrospective measures covering the previous 12 months. [^7] This contrasts with the CPS, where annual outcomes refer to the previous calendar year. These include indicators for employment, usual hours worked per week, number of weeks worked, part-time status, wages and total salary income (earnings). [^8] Throughout the paper, we define intensive margin outcomes such as wages conditional on employment.
WFH Job Postings Data (Hansen et al., 2023)
Our preferred measure of WFH exposure is derived from job postings data on remote and hybrid work arrangements from 2019 to 2024. These data were constructed by a team of researchers in collaboration with Lightcast, using a large language model (LLM) trained to classify job advertisements based on whether the position allows remote work at least one day per week. The resulting measures are available at the month-year–occupation level [^9] Occupations are classified using the Standard Occupational Classification 2018 (SOC) at 3 digits. and at the month-year- industry level. [^10] Industry codes are classified using North American Industry Classification System (NAICS) 2022 at 3 digits. We use this data to build WFH probability changes by occupation, which will be key for our definition of treatment, as we discuss in Subsection 3.3.111111In Section 6, we also use an alternative measure capturing WFH feasibility, that is the teleworkability index à la Dingel and Neiman (2020).
@@
Current Population Survey (CPS) and Survey of Income and Program Participation (SIPP)
We replicate our main findings with public-use microdata from the CPS (Flood et al., 2024) spanning the years 2017–2024, and the 2018-2024 SIPP panel datasets. For the CPS, we integrate data from both the basic monthly files and the Annual Social and Economic Supplement (ASEC), also known as the March files. While weekly measures of employment are available in both the basic monthly and March files, annual employment/labor force participation and income variables are available exclusively in the March files. [^12] See Appendix B of Kleven (2024) for a detailed discussion of differences between the CPS basic monthly and March CPS files, and between weekly and annual variables. In contrast with ACS, annual variables refer to the previous calendar year. In the case of the the SIPP, we exploit the 4-year rotational panel survey. [^13] In 2018 the SIPP redesigned the survey that make it difficult to compare the 2014-2017 SIPP panels and post 2018 panels. The 2018-2024 SIPP panels cover information for reference years 2017-2023. Households are interviewed once a year and asked to recall their detailed monthly employment status and income during the previous year. After the year they join the SIPP sample, households are followed and interviewed for the next three years, after which they are dropped from the sample. Each year new households join the sample, so up to 4 different household cohorts populate each year’s SIPP sample. Importantly, the 2019 SIPP sample experienced significant logistical difficulties, which led the US Census to prioritize resources on following through with the 2018 sample. As a result, the sample of new households that joined in 2019 was not adequately surveyed and the decision was made to not follow up on them in subsequent surveys. Its panel structure allows us to incorporate household fixed effects into our baseline specification to account for time-invariant unobserved heterogeneity. Also, we use SIPP childcare topical modules, which are available for all years in our 2018-2024 SIPP panels, to complement mechanism analysis.